Session: Feb 2011 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (C.Bortolotti at Fri Feb 25 15:39:16 2011) Success:
No known problems

Westerbork (Geert Kuper at Fri Feb 25 16:18:07 2011) Minor failures: FORM
Small syncerror Mark5B formatter from 08:36 UT- 09:34 UT and from 11:09 UT - 11:22 UT

Onsala (Michael Lindqvist at Sun Feb 27 09:33:33 2011) Success:
Apart from RFI, no known problems.

Torun (Kaz Borkowski at Mon Feb 28 08:23:44 2011) Success: POIN
Telescope movements were a bit slower due to cold temperature, so occasionally could be late on source. Otherwise no known problems.

Effelsberg (Uwe Bach at Tue Mar 01 14:30:52 2011) Success:

Jodrell_Bank (Alastair Gunn at Tue Mar 29 11:26:09 2011) Success:
No known problems.